# CVOCSemRPL: Class-Variance Optimized Clustering, Semantic Information Injection and Restricted Pseudo Labeling based Improved Semi-Supervised Few-Shot Learning

### v2

@@ 7.1 Datasets
The miniImageNet dataset [29] is a subset of ImageNet [31], consisting of 60,000 images, from 64 training classes, 16 validation classes, and 20 testing classes. The tieredImageNet dataset [28], is another subset of ImageNet [31] that is designed specifically for few-shot learning applications. This dataset of 779,165 images, includes a hierarchical structure with 34 superclasses divided into 608 fine-grained categories, encompassing a wide range of natural and artificial objects, animal species, and other general image classes. The dataset is partitioned into 20 training superclasses (covering 351 classes), 6 validation superclasses (97 classes), and 8 test superclasses (160 classes). The CUB dataset [30], derived from the fine-grained Caltech-UCSD Birds-200-2011 dataset [32], contains 200 classes representing 11,788 images of birds. This dataset is partitioned into 100, 50 and 50 classes for the base, validation, and novel class split.
7.2 Implementation Details
7.2.1 Pretraining
@@
During the pre-training stage, the CVOCSemRPL model is trained for 200 epochs using a batch size of 128, utilizing all available training categories and data. A dropout rate of 0.1, weight decay of 0.0005, and momentum of 0.9 are applied. The model is optimized using the stochastic gradient descent (SGD) algorithm with an initial learning rate set to 0.1. If the validation loss does not improve for 10 consecutive epochs, the learning rate is reduced by a factor of 10. The output dimensionality of the feature representations depends on the backbone: it is 512 for ResNet-12 and 640 for WRN-28-10.
7.2.2 Finetuning
In the fine-tuning stage, the CVOCSemRPL model is trained for 200 epochs over 600 iterations. The optimization employs stochastic gradient descent (SGD) with a momentum of 0.9 and a weight decay of 0.0005. The learning rate is initially set to 0.001, and it is reduced by a factor of 10 when the validation performance plateaus. Episodic training is done at this stage.
7.2.3 Semantic Injection Network Training
@@
The Semantic Injection Network is trained separately after episodic fine-tuning. We train for 100 epochs with a batch size of 128 support–query pairs drawn from all training classes. The input–output dimensions are 512 for image features when using ResNet-12, and 640 when using WRN-28-10 as the backbone. In both cases, the text features (semantic embeddings from CLIP)( [27]) have a dimensionality of 512. A hidden layer of size 4096 is used. Training is performed using the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, weight decay of $1\times 10^{-4}$, and a dropout rate of 0.1. The learning rate is decayed by a factor of 0.1 every 30 epochs. The total loss is the mean of three L1 components: prototype reconstruction loss, semantic feature reconstruction loss, and joint embedding prototype loss, as detailed in the main paper. The checkpoint with the lowest validation reconstruction loss is selected for downstream few-shot evaluation.
7.2.4 Testing
@@
During the testing phase, the CVOCSemRPL model is evaluated over 1000 episodes, and the mean classification accuracy across these episodes is reported as the final performance metric. Each few-shot task consists of 5 classes (N = 5). For every class, there are K = 1 and K = 5 support samples, along with 15 query samples (q = 15) per class. For each category, the number of unlabeled samples $u$ is 100. The temperature parameter $\tau$ is fixed at $0.1$, and the additional hyperparameter $\varepsilon$ is configured as $0.01$. The clustering process is performed over $10$ iterations.
@@ 7.2.5 Compute Resources
All experiments were conducted on a single NVIDIA A6000 GPU with 48GB of VRAM. The pretraining phases utilized up to 30GB of GPU memory, while the subsequent fine-tuning and testing stages required approximately between 3GB to 8GB.
7.3 Compared Methods
@@
Our comparative experiments include several established baselines as well as our proposed model, CVOCSemRPL. These include cluster-FSL [4], PLML+EP [1], EPNet [3], and ICI [22], along with state-of-the-art few-shot learning methods such as TADAM [11], MTL [2], MetaOpt-SVM [9], CAN [33], LST [34], and LEO [7].
@@
We also evaluate against semi-supervised techniques like TPN [18], TransMatch [13], and PLAIN [35], as well as graph-based models such as wDAE-GNN [8]. Furthermore, comparisons are made with PTN [21] and clustering-based methods introduced by Ren et al. [28], including Soft K-Means, Soft K-Means+Cluster, and Masked Soft K-Means.
@@
Unlike cluster-FSL, which primarily relies on quantitative evaluation of classification accuracy, our proposed method incorporates semantic context, both inter-class and intra-class variance, and unlabeled samples with low prediction entropy to enhance testing performance.
8 More details on CVOC
8.1 Algorithmic Details
8.1.1 Finetuning
@@
In the fine-tuning stage, we invoke Alg. 1, which in each episode computes the reconstruction distance of all support and query embeddings against class-specific dictionaries, applies intra- and inter-class distance regularization and a Cluster Separation Tuner (CST, see Alg. 2) to refine the prototypes, and finally derives query logits (soft labels) from the updated prototypes. Experimentally, CST improves classification accuracy, but introduces additional computational overhead: the algorithm with CST takes approximately 45–46 seconds per epoch (training + validation) on average, whereas without CST it takes only 20–22 seconds per epoch.
@@
Algorithm 1 Class–Variance–Optimised Clustering (CVOC) in the finetuning stage
@@
4:$A\leftarrow S\cup Q$triangleright concatenate support and query sets
@@
21:         $\mu_{c}\leftarrow\frac{1}{|A_{c}|}\sum_{\mathbf{a}\in A_{c}}\mathbf{a}$triangleright mean of current members
@@
23:    $\{\mu_{c}\}\leftarrow\textsc{ClusterSeparationTuner}\bigl{(}\{\mu_{c}\}\bigr{)}$triangleright margin polishing
@@
33:$Q_{\text{cvoc}}\leftarrow\mathrm{softmax}\bigl{(}\ell/\tau\bigr{)},\quad\hat{y}\leftarrow\arg\max_{c}Q_{\text{cvoc}}$
@@
8:             $B_{c}\leftarrow-10^{6}$triangleright $\triangleright$ empty classes stay dim
@@
25:    $\alpha\leftarrow 0.995\,\alpha$triangleright $\triangleright$ anneal random exploration
@@
8.1.2 Testing
At the inference time, we first apply CVOC to the combined support and unlabeled samples to obtain a subset of high-confidence samples; these pseudo-labels are merged into the episode dictionary, and the updated combined support, unlabeled and query embeddings are passed to a Label Propagation module [24], which produces query logits and final predictions.
While semantic anchor module significantly improves accuracy, it has one drawback- computational overhead (see Table 6). For this module, we also introduce an extra hyperparameter that controls the mixing between the visual prototype and its semantic reconstruction (See Section 9.2). Apart from inference, we include a separate step to train the Semantic Injection Network, which takes approximately 30–35 seconds per epoch (training + validation) on average.
@@ 8.1.2 Testing
With semantic anchor 0.330 s 0.355 s
8.2 Why Label Propagation
@@
8.3 Class Description Generation Pipeline
To supply the semantic-anchor module with concise yet visually informative text, we automatically build a short description for every class name [^1] WordNet glosses for each classes are automatically retrieved via the NLTK  [36] interface to WordNet [37]. by running a four–stage sequential agentic chain (see Figure 4) using the GPT-4o-mini model accessed via the OpenAI API [38]. Each stage receives the output of the previous one and is instructed with a fixed prompt; no manual intervention is required.
@@ 8.3 Class Description Generation Pipeline
Table 8: Semi-supervised few-shot classification on miniImageNet dataset ($5$-way 1-shot, ResNet-12 backbone). ”Num_U” represents the number of unlabeled samples.
Num_U Acc (Ours) Cluster-FSL [4]
@@
Table 9: Semi-supervised few-shot classification on miniImageNet dataset ($5$-way 5-shot, ResNet-12 backbone).”Num_U” represents the number of unlabeled samples.
Num_U Acc (Ours) Cluster-FSL [4]
@@
Figure 5: Accuracy comparison across different methods on the miniImageNet dataset using the ResNet-12 backbone for 5-way 1-shot and 5-way 5-shot scenarios. NN: nearest neighbour approach.
@@ 8.3.1 Effectiveness of our approach
Unfortunately, WordNet glosses are short (often $<\!20$ words) and rarely mention fine-grained cues such as colour patches, plumage texture, or limb shape that are crucial for few-shot recognition. We therefore prompt GPT-4o-mini to expand each gloss, injecting the missing visual details.
@@
Iterative self-refinement has been shown to improve factuality and task performance in LLMs [39]. So, our four-stage pipeline writer $\rightarrow$ Filter $\rightarrow$ Visualizer $\rightarrow$ Finalizer realises this idea: the Filter runs at low temperature to deterministically prune non-visual content, while the Visualizer is sampled with a slightly higher temperature to fill in plausible but commonly known traits; the Finalizer then consolidates the description into 50–70 highly informative words. Controlling generation temperature is a well-studied way of trading off creativity and determinism in LLMs.
@@
8.4 Effect of unlabeled samples

### v3

@@ 8.3.1 Datasets
The miniImageNet dataset introduced by vinyals2016matching is a subset of ImageNet by russakovsky2015imagenet, consisting of 60,000 images from 64 training classes, 16 validation classes, and 20 testing classes. The tieredImageNet dataset proposed by ren2018meta is another ImageNet subset designed specifically for few-shot learning applications. This dataset of 779,165 images includes a hierarchical structure with 34 superclasses divided into 608 fine-grained categories, encompassing a wide range of natural and artificial objects, animal species, and other general image classes. It is partitioned into 20 training superclasses (351 classes), 6 validation superclasses (97 classes), and 8 test superclasses (160 classes). The CUB dataset introduced by hilliard2018few, derived from the fine-grained Caltech-UCSD Birds-200-2011 dataset by welinder2010caltech, contains 200 classes representing 11,788 bird images. It is partitioned into 100, 50, and 50 classes for the base, validation, and novel splits.
For the domain adaptation part, we used the following datasets along with the CUB dataset [hilliard2018few] -
  • EuroSAT [helber2019eurosat] is a set of satellite images of landscapes.
  • ISIC [codella2019isic] is a set of dermoscopy images of human skin lesions.
  • ChestX [wang2017chestx] is a set of X-Ray images of the human chest.
Our dataset compilations and splits follow those used by [oh2022understanding]. Figure 6 shows several examples from the target domain datasets.
(a) EuroSAT
(b) ISIC
(c) ChestX
Figure 6: Examples from our target domain datasets for cross-domain experiments.
8.3.2 Implementation Details
Pretraining
@@ 8.3.2 Implementation Details
During the pre-training stage in our approach, the model is trained for 200 epochs using a batch size of 128, utilizing all available training categories and data. A dropout rate of 0.1, weight decay of 0.0005, and momentum of 0.9 are applied. The model is optimized using the stochastic gradient descent (SGD) algorithm with an initial learning rate set to 0.1. If the validation loss does not improve for 10 consecutive epochs, the learning rate is reduced by a factor of 10. The output dimensionality of the feature representations depends on the backbone: it is 512 for ResNet-12 and 640 for WRN-28-10.
Finetuning
In the fine-tuning stage in our approach, the model is trained for 200 epochs over 600 iterations. The optimization employs stochastic gradient descent (SGD) with a momentum of 0.9 and a weight decay of 0.0005. The learning rate is initially set to 0.001, and it is reduced by a factor of 10 when the validation performance plateaus. Episodic training is done at this stage. To keep the additional computational cost minimal and to ensure that CST (see Alg. 2) does not dominate the main CVOC (see Alg. 1) algorithm, we restrict CST to only one iteration. Meanwhile, the number of iterations in CVOC is set to 10 in all settings. The remaining CST hyperparameters used are $\epsilon=2.0$, $\beta_{0}=0.05$, $\gamma=0.005$, $\alpha=0.02$ across every settings and datasets. This conservative choice of CST update parameters ensures that the overall CVOC algorithm, when coupled with CST, remains stable and efficient, making it well-suited for practical deployment scenarios while being robust and not overly sensitive to a high number of hyperparameters. Refer to  8.5.3 to find various ablation studies with $w_{\text{intra}}$ and $w_{\text{inter}}$ and other CST hyperparameters.
Semantic Injection Network Training
@@
The Semantic Injection Network in our approach is trained separately after episodic fine-tuning. We train for 100 epochs with a batch size of 128 support–query pairs drawn from all training classes. The input–output dimensions are 512 for image features when using ResNet-12 and 640 when using WRN-28-10 as the backbone. In both cases, the text features (semantic embeddings from CLIP)( [radford2021learning]) have a dimensionality of 512. A hidden layer of size 4096 is used. Training is performed using the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, a weight decay of $1\times 10^{-4}$, and a dropout rate of 0.1. The learning rate is decayed by a factor of 0.1 every 30 epochs. The total loss is the mean of three L1 components: prototype reconstruction loss, semantic feature reconstruction loss, and joint embedding prototype loss, as detailed in the main text. The checkpoint with the lowest validation reconstruction loss is selected for downstream few-shot evaluation.
Testing
@@
During the testing phase, the model is evaluated over 1000 episodes, and the mean classification accuracy across these episodes is reported as the final performance metric. Each few-shot task consists of 5 classes (N = 5). For every class, there are K = 1 and K = 5 support samples, along with 15 query samples (q = 15) per class. For each category, the number of unlabeled samples $u$ is 100. The temperature parameter $\tau$ is fixed at $0.1$. The clustering process is performed over $10$ iterations.
@@
All experiments were conducted on a single NVIDIA A6000 GPU. The pretraining phase utilized up to 30 GB of GPU memory, while the subsequent fine-tuning and testing stages required approximately 3–6 GB, which is comparable to previously existing methods.
8.3.3 Compared Methods
@@
Our comparative experiments include several established baselines as well as our proposed approach. These include cluster-FSL [ling2022semi], PLML+EP [dong2024pseudo], EPNet [rodriguez2020embedding], and ICI [wang2020instance], along with state-of-the-art few-shot learning methods such as TADAM [oreshkin2018tadam], MTL [sun2019meta], MetaOpt-SVM [lee2019meta], CAN [hou2019cross], LST [sung2018learning], and LEO [rusu2018meta].
@@
We also evaluate against semi-supervised techniques like TPN [liu2018learning], TransMatch [yu2020transmatch], and PLAIN [li2021semi], as well as graph-based models such as wDAE-GNN [gidaris2019generating]. Furthermore, comparisons are made with PTN [huang2021ptn] and clustering-based methods introduced by [ren2018meta], including Soft K-Means, Soft K-Means+Cluster, and Masked Soft K-Means, iLPC [lazarou2025exploiting].
@@
Unlike cluster-FSL, we use both inter-class and intra-class variance, and unlabeled samples with low prediction entropy to enhance testing performance.
8.4 More details on CVOC
8.4.1 Background on Propagation Methods and Inference Paradigms
In semi-supervised few-shot learning (SSFSL), leveraging unlabeled data in addition to very few labeled support examples is a promising way to improve generalization. Two graph-based propagation techniques that have been used in prior works are Embedding Propagation (EP) and Label Propagation (LP). These techniques are especially effective in transductive settings, where the model sees unlabeled query points during adaptation. Below we briefly review their principles, how they have been applied in SSFSL, and then state our use of them.
Embedding Propagation (EP)
Embedding Propagation was introduced by [rodriguez2020embedding], which enforces smoothness over the embedding space by interpolating embeddings along a similarity graph. After extracting features for all samples, it constructs a similarity graph and propagates embeddings through this graph to obtain smoothed embeddings. This smoothing helps reduce noise and makes decision boundaries more robust, which is particularly useful in semi-supervised or transductive few-shot settings.
Label Propagation (LP)
Label Propagation was introduced by zhu2002lp as a classical method in which a small number of nodes have known labels, and labels are diffused through a graph over unlabeled nodes. In few-shot and SSFSL domains, LP has been widely used to propagate label information from support examples to query or unlabeled points, improving classification performance in low-label regimes. LP has been adopted in several notable SSFSL methods, including [liu2018learning] and other graph-based SSFSL approaches that refine or expand label information through unlabeled samples.
Usage in SSFSL and Prior Works
Across the SSFSL literature, EP and LP have become standard building blocks. EP has been used to regularize embedding spaces, while LP has been employed to propagate supervision to unlabeled data. Methods such as EPNet, TPN, and related propagation-based approaches highlight the central role of these techniques in advancing SSFSL.
Our Usage
In our method, we also adopt both EP and LP. We employ EP to smooth embeddings over the graph of support, query, and unlabeled data, and then apply LP to infer labels for unlabeled and query samples. This design choice aligns our approach with prior SSFSL works while contributing to improved generalization.
8.4.2 Algorithmic Details
Finetuning
@@ 8.4.2 Algorithmic Details
In the fine-tuning stage, we invoke Alg. 1, which in each episode computes the reconstruction distance of all support and query embeddings against class-specific dictionaries, applies intra- and inter-class distance regularization and a Cluster Separation Tuner (CST, see Alg. 2) to refine the prototypes, and finally derives query logits (soft labels) from the updated prototypes. Experimentally, CST improves classification accuracy, but introduces additional computational overhead: the algorithm with CST takes approximately 10-11 seconds per epoch on average, whereas without CST it takes only 8-9 seconds per epoch. We found that CST works best with low iterations (1-3). Higher iterations (5-10) can cause prototypes to drift too far, leading to instability, and also increase the training time too much (with 10 iterations, training time per epoch becomes 1.8x). Also, the "1 iteration" choice is a deliberate design for Efficiency. Since the prototypes are already initialized near the true mean by CVOC, a single CST step acts as a fine-grained "boundary sharpening" operation rather than a global optimization, ensuring stability and minimal computational cost.
@@
Algorithm 1 Class–Variance–Optimized Clustering (CVOC) in the finetuning stage
@@
[…]
@@
21:    $\mu_{c}\leftarrow\frac{1}{|A_{c}|}\sum_{\mathbf{a}\in A_{c}}\mathbf{a}$ $\triangleright$ $\triangleright$ mean of current members
@@
23:  $\{\mu_{c}\}\leftarrow\textsc{ClusterSeparationTuner}\bigl(\{\mu_{c}\}\bigr)$ $\triangleright$ $\triangleright$ margin polishing
@@
[…]
@@
8:     $B_{c}\leftarrow-10^{6}$ $\triangleright$ $\triangleright$ empty classes stay dim
@@
25:  $\alpha\leftarrow 0.995\,\alpha$ $\triangleright$ $\triangleright$ anneal random exploration
@@
Testing
At the inference time, we first apply CVOC to the combined support and unlabeled samples to obtain a subset of high-confidence samples and their pseudo-labels. These pseudo-labels are merged into the episode dictionary, and the updated combined support, unlabeled, and query embeddings are passed to a Label Propagation module [iscen2019label], which produces query logits and final predictions.
The semantic anchor module significantly improves accuracy, while only slightly increasing the inference time, as can be seen in Table 9. For this module, we also introduce an extra hyperparameter that controls the mixing between the visual prototype and its semantic reconstruction (See Section 8.5.5). Before inference, we include a separate step to train the Semantic Injection Network, which takes approximately 10-12 seconds per epoch on average.
@@
With semantic anchor 0.187 s 0.208 s
8.4.3 Why Label Propagation
@@
8.4.4 Class Description Generation Pipeline
To supply the semantic‐anchor module with concise yet visually informative text, we automatically build a short description for every class name [^1] WordNet glosses for each classes are automatically retrieved via the NLTK  [bird2009nltk] interface to WordNet [fellbaum1998wordnet]. by running a four–stage sequential agentic chain (see Figure 7) using the GPT-4o-mini model accessed via the OpenAI API [openai_gpt4o_2024]. Each stage receives the output of the previous one and is instructed with a fixed prompt; no manual intervention is required.
@@ 8.4.4 Class Description Generation Pipeline
Table 11: Ablation Study on Description Generation Strategy for Semantic Anchoring. We provide accuracy on the miniImageNet dataset for the 5-way 1-shot task, using ViT-B/32 CLIP encoding and a ResNet-12 classifier backbone.
Description Strategy Description Source Accuracy (%)
(miniImageNet 5-way 1-shot)
Baseline Class Name Only 78.33
Raw Text Input Unedited WordNet Gloss 80.81
LLM Baseline Single-Shot Prompt 82.55
Our Proposed Method Agentic Chain (4-Stage Pipeline) 84.51
@@
Table 12: Semi-supervised few-shot classification on miniImageNet dataset (55-way 1-shot, ResNet-12 backbone). ”Num_U” represents the number of unlabeled samples.
Num_U Acc (Ours) Cluster-FSL [ling2022semi]
@@
Table 13: Semi-supervised few-shot classification on miniImageNet dataset (55-way 5-shot, ResNet-12 backbone).”Num_U” represents the number of unlabeled samples.
Num_U Acc (Ours) Cluster-FSL [ling2022semi]
@@
Figure 8: Accuracy comparison across different methods on the miniImageNet dataset using the ResNet-12 backbone for 5-way 1-shot and 5-way 5-shot scenarios. NN: nearest neighbour approach. SIN: Semantic Injection. C-FSL: Cluster-FSL.
@@
Unfortunately, WordNet glosses are short (often <20<\!20 words) and rarely mention fine-grained cues such as colour patches, plumage texture, or limb shape that are crucial for few-shot recognition. We therefore prompt GPT-4o-mini to expand each gloss, injecting the missing visual details.
@@
Iterative self-refinement has been shown to improve factuality and task performance in LLMs [madaan2023selfrefine]. So, our four-stage pipeline Writer $\rightarrow$ Filter $\rightarrow$ Visualizer $\rightarrow$ Finalizer realises this idea: the Filter runs at low temperature to deterministically prune non-visual content, while the Visualizer is sampled with a slightly higher temperature to fill in plausible but commonly known traits; the Finalizer then consolidates the description into 50–70 highly informative words. Controlling generation temperature is a well-studied way of trading off creativity and determinism in LLMs.
@@
Agentic Chain Sensitivity Analysis
To further validate the robustness of our agentic architecture, we conducted a sensitivity analysis on various CLIP models and LLMs within our framework. Our primary experiments were conducted using the ViT-B/32 CLIP model. However, as demonstrated in Figure 10(b), we observe remarkably similar performance across different ViT architectures, including ViT-B/32, ViT-B/16, and ViT-L/14. This suggests that the benefits of our agentic chain are not solely dependent on the most powerful visual encoder, as even with lower-version CLIP models, we achieve comparable results. Specifically, on the miniImageNet dataset with a 5-way 1-shot task using a ResNet-12 backbone, we obtained an accuracy of 84.51% with ViT-B/32, 84.52% with ViT-B/16, and 86.63% with ViT-L/14. This indicates that our architecture effectively leverages the visual information regardless of the precise CLIP model employed.
Furthermore, we investigated the impact of different LLMs on the performance of our agentic chain. As shown in Figure 10(a), our architecture demonstrates consistent strong performance even when integrating weaker LLMs. We hypothesize that the agentic nature of our framework, which facilitates iterative refinement and contextual reasoning, allows for the generation of superior visual description-based texts, thereby mitigating the performance limitations typically associated with less powerful language models. For instance, with GPT-4o-mini, we achieved an accuracy of 84.51%, closely followed by LLama 3.1 8B [touvron2023llama] at 84.31%, Qwen 2.5 7B [qwen2_5_2024] at 84.6%, and Qwen 2.5 3B [qwen2_5_2024] at 84.40%. These results underscore the potential of our agentic architecture to democratize access to high-quality visual understanding, enabling effective performance even with computationally lighter LLMs. All LLM experiments were also conducted on the miniImageNet 5-way 1-shot task with a ResNet-12 backbone.
(a) LLM Sensitivity
(b) CLIP Sensitivity
Figure 10: Sensitivity Analysis of the Agentic Architecture. (a) Performance variations across different Large Language Models (LLMs). The plot demonstrates that our agentic framework maintains robust accuracy even when utilizing smaller, computationally efficient models (e.g., Qwen 2.5 3B), supporting the hypothesis that iterative agentic reasoning compensates for lower raw LLM capability. (b) Performance variations across different CLIP visual backbones. While larger models like ViT-L/14 yield the highest accuracy, the system achieves highly competitive results with lighter architectures (ViT-B/32, ViT-B/16), indicating that the method effectively extracts visual signals regardless of encoder capacity. All results are reported on the miniImageNet 5-way 1-shot task using a ResNet-12 backbone.
8.4.5 Effect of unlabeled samples